$SPAD/input schaum20.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.429 ∫ tan ax dx

∫ tan ax = −\frac{1}{a} ln cos ax = \frac{1}{a} ln sec ax

— * —

)set break resume
)spool schaum20.output
)set message test on
)set message auto off
)clear all
--S 1 of 56
aa:=integrate(tan(a*x),x)
--R
--R
--R                    2
--R        log(tan(a x)  + 1)
--R   (1)  ------------------
--R                2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 56
bb1:=-1/a*log(cos(a*x))
--R
--R          log(cos(a x))
--R   (2)  - -------------
--R                a
--R                                                    Type: Expression(Integer)
--E
--S 3 of 56
bb2:=1/a*log(sec(a*x))
--R
--R        log(sec(a x))
--R   (3)  -------------
--R              a
--R                                                    Type: Expression(Integer)
--E
--S 4 of 56
cc1:=aa-bb1
--R
--R                    2
--R        log(tan(a x)  + 1) + 2log(cos(a x))
--R   (4)  -----------------------------------
--R                         2a

--R                                                    Type: Expression(Integer)
--E
--S 5 of 56
tanrule:=rule(tan(a) == sin(a)/cos(a))
--R
--R                  sin(a)
--R   (5)  tan(a) == ------
--R                  cos(a)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 6 of 56
dd1:=tanrule cc1
--R
--R                    2           2
--R            sin(a x)  + cos(a x)
--R        log(---------------------) + 2log(cos(a x))
--R                          2
--R                  cos(a x)
--R   (6)  -------------------------------------------
--R                             2a
--R                                                    Type: Expression(Integer)
--E
--S 7 of 56
ee1:=expandLog dd1
--R
--R                    2           2
--R        log(sin(a x)  + cos(a x) )
--R   (7)  --------------------------
--R                    2a
--R                                                    Type: Expression(Integer)
--E
--S 8 of 56
sincossqrrule:=rule(sin(a)^2+cos(a)^2 == 1)
--R
--R              2         2
--I   (8)  sin(a)  + cos(a)  + %K == %K + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 9 of 56      14:429 Schaums and Axiom agree
ff1:=sincossqrrule ee1
--R
--R   (9)  0
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.430 ∫ tan^{2} ax dx

∫ tan^{2} ax = \frac{tan ax}{x} − x

— * —

)clear all
--S 10 of 56
aa:=integrate(tan(a*x)^2,x)
--R
--R
--R        tan(a x) - a x
--R   (1)  --------------
--R               a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 11 of 56
bb:=tan(a*x)/a-x
--R
--R        tan(a x) - a x
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 12 of 56     14:430 Schaums and Axiom agree
cc:=aa-bb
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.431 ∫ tan^{3} ax dx

∫ tan^{3} ax = \frac{tan^{2} ax}{2a} + \frac{1}{a} ln cos ax

— * —

)clear all
--S 13 of 56
aa:=integrate(tan(a*x)^3,x)
--R
--R
--R                      2                2
--R        - log(tan(a x)  + 1) + tan(a x)
--R   (1)  --------------------------------
--R                       2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 14 of 56
bb:=tan(a*x)^2/(2*a)+1/a*log(cos(a*x))
--R
--R                                 2
--R        2log(cos(a x)) + tan(a x)
--R   (2)  --------------------------
--R                    2a
--R                                                    Type: Expression(Integer)
--E
--S 15 of 56
cc:=aa-bb
--R
--R                      2
--R        - log(tan(a x)  + 1) - 2log(cos(a x))
--R   (3)  -------------------------------------
--R                          2a
--R                                                    Type: Expression(Integer)
--E
--S 16 of 56
tanrule:=rule(tan(a) == sin(a)/cos(a))
--R
--R                  sin(a)
--R   (4)  tan(a) == ------
--R                  cos(a)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 17 of 56
dd:=tanrule cc
--R
--R                      2           2
--R              sin(a x)  + cos(a x)
--R        - log(---------------------) - 2log(cos(a x))
--R                            2
--R                    cos(a x)

--R   (5)  ---------------------------------------------
--R                              2a
--R                                                    Type: Expression(Integer)
--E
--S 18 of 56
ee:=expandLog dd
--R
--R                      2           2
--R          log(sin(a x)  + cos(a x) )
--R   (6)  - --------------------------
--R                      2a
--R                                                    Type: Expression(Integer)
--E
--S 19 of 56
sincossqrrule:=rule(sin(a)^2+cos(a)^2 == 1)
--R
--R              2         2
--I   (7)  sin(a)  + cos(a)  + %L == %L + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 20 of 56     14:431 Schaums and Axiom agree
ff:=sincossqrrule ee
--R
--R   (8)  0
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.432 ∫ tan^{n} ax sec^{2} ax dx

∫ tan^{n} ax sec^{2} ax = \frac{tan^{n+1} ax}{(n + 1)a}

— * —

)clear all
--S 21 of 56
aa:=integrate(tan(a*x)^n*sec(a*x)^2,x)
--R
--R
--R                        sin(a x)
--R                  n log(--------)
--R                        cos(a x)
--R        sin(a x)%e

--R   (1)  -------------------------
--R            (a n + a)cos(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 22 of 56
bb:=tan(a*x)^(n+1)/((n+1)*a)
--R
--R                n + 1
--R        tan(a x)
--R   (2)  -------------
--R           a n + a
--R                                                    Type: Expression(Integer)
--E
--S 23 of 56
cc:=aa-bb
--R
--R                        sin(a x)
--R                  n log(--------)
--R                        cos(a x)                    n + 1
--R        sin(a x)%e                - cos(a x)tan(a x)
--R   (3)  -------------------------------------------------
--R                        (a n + a)cos(a x)
--R                                                    Type: Expression(Integer)
--E
--S 24 of 56
explog:=rule(%e^(n*log(x)) == x^n)
--R
--R          n log(x)     n
--R   (4)  %e         == x
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 25 of 56
dd:=explog cc
--R
--R                          n + 1            sin(a x) n
--R        - cos(a x)tan(a x)      + sin(a x)(--------)
--R                                           cos(a x)
--R   (5)  ---------------------------------------------
--R                      (a n + a)cos(a x)
--R                                                    Type: Expression(Integer)
--E
--S 26 of 56
tanrule:=rule(tan(a) == sin(a)/cos(a))
--R
--R                  sin(a)

--R   (6)  tan(a) == ------
--R                  cos(a)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 27 of 56
ee:=tanrule dd
--R
--R                   sin(a x) n + 1            sin(a x) n
--R        - cos(a x)(--------)      + sin(a x)(--------)
--R                   cos(a x)                  cos(a x)
--R   (7)  -----------------------------------------------
--R                       (a n + a)cos(a x)
--R                                                    Type: Expression(Integer)
--E
--S 28 of 56     14:432 Schaums and Axiom agree
ff:=complexNormalize ee
--R
--R   (8)  0
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.433 ∫ \begin{matrix}sec & 2 & ax \\ tan & ax \end{matrix} dx

∫ \begin{matrix}sec & 2 & ax \\ tan & ax \end{matrix} = \frac{1}{a} ln tan ax

— * —

)clear all
--S 29 of 56
aa:=integrate(sec(a*x)^2/tan(a*x),x)
--R
--R
--R              sin(a x)              2cos(a x)
--R        log(------------) - log(- ------------)
--R            cos(a x) + 1          cos(a x) + 1
--R   (1)  ---------------------------------------
--R                           a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 30 of 56
bb:=1/a*log(tan(a*x))

--R
--R        log(tan(a x))
--R   (2)  -------------
--R              a
--R                                                    Type: Expression(Integer)
--E
--S 31 of 56
cc:=aa-bb
--R
--R                                sin(a x)              2cos(a x)
--R        - log(tan(a x)) + log(------------) - log(- ------------)
--R                              cos(a x) + 1          cos(a x) + 1
--R   (3)  ---------------------------------------------------------
--R                                    a
--R                                                    Type: Expression(Integer)
--E
--S 32 of 56
dd:=expandLog cc
--R
--R        - log(tan(a x)) + log(sin(a x)) - log(cos(a x)) - log(- 2)
--R   (4)  ----------------------------------------------------------
--R                                     a
--R                                                    Type: Expression(Integer)
--E
--S 33 of 56     14:433 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R          log(- 2)
--R   (5)  - --------
--R              a
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.434 ∫ \frac{dx}{tan ax} dx

∫ \frac{1}{tan ax} = \frac{1}{a} ln sin ax

— * —

)clear all
--S 34 of 56

aa:=integrate(1/tan(a*x),x)
--R
--R
--R                      2
--R        - log(tan(a x)  + 1) + 2log(tan(a x))
--R   (1)  -------------------------------------
--R                          2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 35 of 56
bb:=1/a*log(sin(a*x))
--R
--R        log(sin(a x))
--R   (2)  -------------
--R              a
--R                                                    Type: Expression(Integer)
--E
--S 36 of 56
cc:=aa-bb
--R
--R                      2
--R        - log(tan(a x)  + 1) + 2log(tan(a x)) - 2log(sin(a x))
--R   (3)  ------------------------------------------------------
--R                                  2a
--R                                                    Type: Expression(Integer)
--E
--S 37 of 56
complexNormalize cc
--R
--R   (4)  0
--R                                                    Type: Expression(Integer)
--E

———-

7 [?]:14.435 ∫ x tan ax dx

∫ x tan ax = \frac{1}{a^{2}} \{\frac{(ax)^{3}}{3} + \frac{(ax)^{5}}{15} + \frac{2(ax)^{7}}{105} + ⋯ + \frac{2^{2n}(2^{2n} − 1)B_{n}(ax)^{2n+1}}{(2n + 1)!} + ⋯\}

— * —

)clear all
--S 38 of 56     14:435 Axiom cannot compute this integral
aa:=integrate(x*tan(a*x),x)

--R
--R
--R           x
--R         ++
--I   (1)   |   %I tan(%I a)d%I
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

8 [?]:14.436 ∫ \frac{tan ax}{x} dx

∫ \frac{tan ax}{x} = ax + \frac{(ax)^{3}}{9} + \frac{2(ax)^{5}}{75} + ⋯ + \frac{2^{2n}(2^{2n} − 1)B_{n}(ax)^{2n−1}}{(2n − 1)(2n)!} + ⋯

— * —

)clear all
--S 39 of 56     14:436 Axiom cannot compute this integral
aa:=integrate(tan(a*x)/x,x)
--R
--R
--R           x
--I         ++  tan(%I a)
--I   (1)   |   --------- d%I
--I        ++       %I
--R                                         Type: Union(Expression(Integer),...)
--E

———-

9 [?]:14.437 ∫ x tan^{2} ax dx

∫ x tan^{2} ax = \frac{x tan ax}{a} + \frac{1}{a^{2}} ln cos ax − \frac{x^{2}}{2}

— * —

)clear all
--S 40 of 56
aa:=integrate(x*tan(a*x)^2,x)
--R
--R
--R                      2                         2 2

--R        - log(tan(a x)  + 1) + 2a x tan(a x) - a x
--R   (1)  -------------------------------------------
--R                              2
--R                            2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 41 of 56
bb:=(x*tan(a*x))/a+1/a^2*log(cos(a*x))-x^2/2
--R
--R                                          2 2
--R        2log(cos(a x)) + 2a x tan(a x) - a x
--R   (2)  -------------------------------------
--R                           2
--R                         2a
--R                                                    Type: Expression(Integer)
--E
--S 42 of 56
cc:=aa-bb
--R
--R                      2
--R        - log(tan(a x)  + 1) - 2log(cos(a x))
--R   (3)  -------------------------------------
--R                           2
--R                         2a
--R                                                    Type: Expression(Integer)
--E
--S 43 of 56
tanrule:=rule(tan(a) == sin(a)/cos(a))
--R
--R                  sin(a)
--R   (4)  tan(a) == ------
--R                  cos(a)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 44 of 56
dd:=tanrule cc
--R
--R                      2           2
--R              sin(a x)  + cos(a x)
--R        - log(---------------------) - 2log(cos(a x))
--R                            2
--R                    cos(a x)
--R   (5)  ---------------------------------------------
--R                               2
--R                             2a
--R                                                    Type: Expression(Integer)

--E
--S 45 of 56
ee:=expandLog dd
--R
--R                      2           2
--R          log(sin(a x)  + cos(a x) )
--R   (6)  - --------------------------
--R                        2
--R                      2a
--R                                                    Type: Expression(Integer)
--E
--S 46 of 56
sincossqrrule:=rule(sin(a)^2+cos(a)^2 == 1)
--R
--R              2         2
--I   (7)  sin(a)  + cos(a)  + %R == %R + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 47 of 56     14:437 Schaums and Axiom agree
ff:=sincossqrrule ee
--R
--R   (8)  0
--R                                                    Type: Expression(Integer)
--E

———-

10 [?]:14.438 ∫ \frac{dx}{p + q tan ax} dx

∫ \frac{1}{p + q tan ax} = \frac{px}{p^{2} + q^{2}} + \frac{q}{a(p^{2} + q^{2})} ln(q sin ax + p cos ax)

— * —

)clear all
--S 48 of 56
aa:=integrate(1/(p+q*tan(a*x)),x)
--R
--R
--R                        2
--R        - q log(tan(a x)  + 1) + 2q log(q tan(a x) + p) + 2a p x
--R   (1)  --------------------------------------------------------
--R                                  2       2
--R                              2a q  + 2a p

--R                                         Type: Union(Expression(Integer),...)
--E
--S 49 of 56
bb:=(p*x)/(p^2+q^2)+q/(a*(p^2+q^2))*log(q*sin(a*x)+p*cos(a*x))
--R
--R        q log(q sin(a x) + p cos(a x)) + a p x
--R   (2)  --------------------------------------
--R                         2      2
--R                      a q  + a p
--R                                                    Type: Expression(Integer)
--E
--S 50 of 56
cc:=aa-bb
--R
--R   (3)
--R                       2
--R       - q log(tan(a x)  + 1) + 2q log(q tan(a x) + p)
--R     +
--R       - 2q log(q sin(a x) + p cos(a x))
--R  /
--R         2       2
--R     2a q  + 2a p
--R                                                    Type: Expression(Integer)
--E
--S 51 of 56
tanrule:=rule(tan(a) == sin(a)/cos(a))
--R
--R                  sin(a)
--R   (4)  tan(a) == ------
--R                  cos(a)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 52 of 56
dd:=tanrule cc
--R
--R   (5)
--R                       2           2
--R               sin(a x)  + cos(a x)
--R       - q log(---------------------) - 2q log(q sin(a x) + p cos(a x))
--R                             2
--R                     cos(a x)
--R     +
--R              q sin(a x) + p cos(a x)
--R       2q log(-----------------------)
--R                      cos(a x)
--R  /

--R         2       2
--R     2a q  + 2a p
--R                                                    Type: Expression(Integer)
--E
--S 53 of 56
ee:=expandLog dd
--R
--R                        2           2
--R          q log(sin(a x)  + cos(a x) )
--R   (6)  - ----------------------------
--R                      2       2
--R                  2a q  + 2a p
--R                                                    Type: Expression(Integer)
--E
--S 54 of 56
sincossqrrule:=rule(sin(a)^2+cos(a)^2 == 1)
--R
--R              2         2
--I   (7)  sin(a)  + cos(a)  + %S == %S + 1
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 55 of 56     14:438 Schaums and Axiom agree
ff:=sincossqrrule ee
--R
--R   (8)  0
--R                                                    Type: Expression(Integer)
--E

———-

11 [?]:14.439 ∫ tan^{n} ax dx

∫ tan^{n} ax = \frac{tan^{n−1} ax}{(n − 1)a} − ∫ tan^{n−2} ax

— * —

)clear all
--S 56 of 56     14:439 Axiom cannot compute this integral
aa:=integrate(tan(a*x)^n,x)
--R
--R
--R           x
--R         ++           n
--I   (1)   |   tan(%I a) d%I

--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 p80